OpenStudy (anonymous):

When solving the system x + 2y = 9 and x - y = 6, which expression could be substituted for x in the second equation? 9 - x y - 9/2 9 - 2y 1/2 y + 9

OpenStudy (across):

x + 2y = 9 x = 9 - 2y

OpenStudy (lilg132):

x = 9 - 2y